The population of a village was 20,000. Within 2 years, the population is increased 3% by birth and 2% by immigration. What will be the population of the village after 2 years? Find it.

Grade:10

2 Answers

Arun
4 years ago
Dear student
 
total population after 2 years = 20000 + 20000*3/100 + 20000* 2/100 = 21000
 
Regards
Arun
Vikas TU
4 years ago
Dear student 
Population = 20000, 
Population increase by Birth = 3% 
Populaton increase by immigraton = 2% 
Total increase in the population after 2 year = 5%
Population increase = 20000 + 20000*5/100 = 21000